OpenStudy (anonymous):

An angle measuring pi/8 radians is equal to which of the angle measures given below? Check all that apply.

OpenStudy (anonymous):

A. 22.5 degrees B. 22 degrees 30' C. 22 degrees 5' D. 39.25 degrees

OpenStudy (solomonzelman):

\(\large\color{black}{\pi=180^\circ }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE \pi }{\LARGE 8}~~~\Rightarrow~~~\frac{\LARGE 180^\circ }{\LARGE 8} ~~~=?}\)

OpenStudy (solomonzelman):

if you are confused, please don't leave like many other newbies. Rather, reply... we will address all math and other site-related and non-related concerns.